In a survey of 2655 ​adults, 1430 say they have started paying bills online in the last year. Construct a​ 99% confidence interval for the population proportion. Interpret the results. A​ 99% confidence interval for the population proportion is
Answer #1

Solution

given 1430, n 2655

1430 sample proportion (p) ; 0.539 2655

for 99% cofidence level, Za 2.576

formula : con fidence interval for the population proportion, p

p* (1 - p)

0.539 (1 0.539) 0.5392.576 2655

0.539 0.461 0.5392.576 2655

0.5390.025

lower bound = 0.539-0.025 0.514

upper bound = 0.539 + 0.025 = 0.564

A​ 99% confidence interval for the population proportion is 0.514, 0.564

With 99% confidence, it can be said that the population proportion of adults who say they have started paying bills online in the last year is between the endpoints of the given confidence interval.